Comparative financial statements for Weller Corporation, a merchandising company, for the year ending December 31 appear below. The company did not issue any new common stock during the year. A total of 910,000 shares of common stock were outstanding. The interest rate on the bond payable was 12%, the income tax rate was 40%, and the dividend per share of common stock was $0.75 last year and $0.40 this year. The market value of the company’s common stock at the end of the year was $23.50. All of the company’s sales are on account.

Weller Corporation
Comparative Balance Sheet
(dollars in thousands)
This Year Last Year
Assets
Current assets:
Cash $ 1,506 $ 1,720
Accounts receivable, net 15,500 10,300
Inventory 10,250 8,640
Prepaid expenses 1,910 2,320
Total current assets 29,166 22,980
Property and equipment:
Land 7,100 7,100
Buildings and equipment, net 20,300 20,100
Total property and equipment 27,400 27,200
Total assets $ 56,566 $ 50,180
Liabilities and Stockholders' Equity
Current liabilities:
Accounts payable $ 10,600 $ 8,850
Accrued liabilities 820 1,250
Notes payable, short term 410 410
Total current liabilities 11,830 10,510
Long-term liabilities:
Bonds payable 6,250 6,250
Total liabilities 18,080 16,760
Stockholders' equity:
Common stock 910 910
Additional paid-in capital 4,750 4,750
Total paid-in capital 5,660 5,660
Retained earnings 32,826 27,760
Total stockholders' equity 38,486 33,420
Total liabilities and stockholders' equity $ 56,566 $ 50,180
Weller Corporation
Comparative Income Statement and Reconciliation
(dollars in thousands)
This Year Last Year
Sales $ 90,000 $ 85,000
Cost of goods sold 57,500 53,500
Gross margin 32,500 31,500
Selling and administrative expenses:
Selling expenses 9,600 9,100
Administrative expenses 13,100 12,100
Total selling and administrative expenses 22,700 21,200
Net operating income 9,800 10,300
Interest expense 750 750
Net income before taxes 9,050 9,550
Income taxes 3,620 3,820
Net income 5,430 5,730
Dividends to common stockholders 364 728
Net income added to retained earnings 5,066 5,002
Beginning retained earnings 27,760 22,758
Ending retained earnings $ 32,826 $ 27,760

Required:

Compute the following financial data for this year:

1. Earnings per share. (Round your answer to 2 decimal places.)

2. Price-earnings ratio. (Round your intermediate calculations and final answer to 2 decimal places.)

3. Dividend payout ratio. (Round your intermediate calculations and final answer to 2 decimal places.)

4. Dividend yield ratio. (Round your answer to 2 decimal places.)

5. Book value per share. (Round your answer to 2 decimal places.)
Answer #1
1. Earnings per share = ( Net income - Preferred dividend ) / Outstanding shares of common stock = ( 5430000 - 0 ) / 910000 5.97
2. Price-earnings ratio = Market price per share / Earnings per share = 23.50 / 5.97 3.94
3. Dividend payout ratio = Dividend per share / Earnings per share = 0.40 / 5.97 6.70%
4. Dividend yield ratio = Dividend per share / Market price per share = 0.40 / 23.50 1.70%
5. Book value per share = ( Total stockholder's equity - Preferred equity ) / Outstanding shares of common stock = ( 38486000 - 0 ) / 910000 42.29